Solve the following national income models by Cramer’s rule Y=C+I+1000,C=10+0.7(Y-T),I=100+0.2Y,T=0.3Y Where Y , C , I, M and T are national income, consumption, investment imports and taxes.
1
Expert's answer
2020-05-20T19:32:57-0400
"\begin{alignedat}{2}
 C+ I-Y=-1000 \\
 C+0.7T-0.7Y =10 \\
 I-0.2Y =100 \\
 T-0.3Y =0
\end{alignedat}"

"\varDelta=\begin{vmatrix}
 1 & 1 & 0 & -1 \\
 1 & 0 & 0.7 & -0.7 \\
 0 & 1 & 0 & -0.2 \\
 0 & 0 & 1 & -0.3 
\end{vmatrix}=""=\begin{vmatrix}
 0 & 0.7 & -0.7 \\
 1 & 0 & -0.2 \\
 0 & 1 & -0.3
\end{vmatrix}-\begin{vmatrix}
 1 & 0 & -1 \\
 1 & 0 & -0.2 \\
 0 & 1 & -0.3
\end{vmatrix}=""=
-\begin{vmatrix}
 0.7 & -0.7 \\
 1 & -0.3
\end{vmatrix}-\begin{vmatrix}
 0 & -0.2 \\
 1 & -0.3
\end{vmatrix}+\begin{vmatrix}
 0 & -1 \\
 1 & -0.3
\end{vmatrix}=""=0.21-0.7-0.2+1=0.31\not=0"

"\varDelta_C=\begin{vmatrix}
 -1000 & 1 & 0 & -1 \\
 10 & 0 & 0.7 & -0.7 \\
 100 & 1 & 0 & -0.2 \\
 0 & 0 & 1 & -0.3 
\end{vmatrix}=""=-\begin{vmatrix}
 10 & 0.7 & -0.7 \\
 100 & 0 & -0.2 \\
 0 & 1 & -0.3
\end{vmatrix}-\begin{vmatrix}
 -1000 & 0 & -1 \\
 10 & 0.7 & -0.7 \\
 0 & 1 & -0.3
\end{vmatrix}=""=
-10\begin{vmatrix}
 0 & -0.2 \\
 1 & -0.3
\end{vmatrix}+100\begin{vmatrix}
 0.7 & -0.7 \\
 1 & -0.3
\end{vmatrix}-""-0.7\begin{vmatrix}
 -1000 & -1 \\
 0 & -0.3
\end{vmatrix}+\begin{vmatrix}
 -1000 & -1 \\
 10 & -0.7
\end{vmatrix}=""=-2-21+70-210+700+10=547"

"\varDelta_I=\begin{vmatrix}
 1 & -1000 & 0 & -1 \\
 1 & 10 & 0.7 & -0.7 \\
 0 & 100 & 0 & -0.2 \\
 0 & 0 & 1 & -0.3 
\end{vmatrix}=""=\begin{vmatrix}
 10 & 0.7 & -0.7 \\
 100 & 0 & -0.2 \\
 0 & 1 & -0.3
\end{vmatrix}-\begin{vmatrix}
 -1000 & 0 & -1 \\
 100 & 0 & -0.2 \\
 0 & 1 & -0.3
\end{vmatrix}=""=
10\begin{vmatrix}
 0 & -0.2 \\
 1 & -0.3
\end{vmatrix}-100\begin{vmatrix}
 0.7 & -0.7 \\
 1 & -0.3
\end{vmatrix}+""+1000\begin{vmatrix}
 0 & -0.2 \\
 1 & -0.3
\end{vmatrix}+100\begin{vmatrix}
 0 & -1 \\
 1 & -0.3
\end{vmatrix}=""=2+21-70+200+100=253"

"\varDelta_T=\begin{vmatrix}
 1 & 1 & -1000 & -1 \\
 1 & 0 & 10 & -0.7 \\
 0 & 1 & 100 & -0.2 \\
 0 & 0 & 0 & -0.3 
\end{vmatrix}=""=-0.3\begin{vmatrix}
 1 & 1 & -1000 \\
 1 & 0 & 10 \\
 0 & 1 & 100
\end{vmatrix}=""=
-0.3\begin{vmatrix}
 0 & 10 \\
 1 & 100
\end{vmatrix}+0.3\begin{vmatrix}
 1 & -1000 \\
 1 & 100
\end{vmatrix}=""=3+30+300=333"

"\varDelta_Y=\begin{vmatrix}
 1 & 1 & 0 & -1000 \\
 1 & 0 & 0.7 & 10 \\
 0 & 1 & 0 & 100 \\
 0 & 0 & 1 & 0 
\end{vmatrix}=""=-\begin{vmatrix}
 1 & 1 & -1000 \\
 1 & 0 & 10 \\
 0 & 1 & 100
\end{vmatrix}=""=
-\begin{vmatrix}
 0 & 10 \\
 1 & 100
\end{vmatrix}+\begin{vmatrix}
 1 & -1000 \\
 1 & 100
\end{vmatrix}=""=10+100+1000=1110"


"C={\varDelta_C\over \varDelta}={547\over0.31}={54700\over31}"

"I={\varDelta_I\over \varDelta}={253\over0.31}={25300\over31}"

"T={\varDelta_T\over \varDelta}={333\over0.31}={33300\over31}"

"Y={\varDelta_Y\over \varDelta}={1110\over0.31}={111000\over31}"